# Flexible inference of learning rules from de novo learning data using neural networks

### v1

@@ Abstract
Understanding how animals learn is a central challenge in neuroscience, with growing relevance to the development of animal- or human-aligned artificial intelligence. However, most existing approaches assume specific parametric forms for the learning rule (e.g., Q-learning, policy gradient) or are limited to simplified settings like bandit tasks, which do not involve learning a new input–output mapping from scratch. In contrast, animals must often learn new behaviors de novo, which poses a rich challenge for learning-rule inference. We target this problem by inferring learning rules directly from animal decision-making data during de novo task learning, a setting that requires models flexible enough to capture suboptimality, history-dependence, and rich external stimulus integration without strong structural priors. We first propose a nonparametric framework that parameterizes the per-trial update of policy weights with a deep neural network (DNN), and validate it by recovering ground-truth rules in simulation. We then extend to a recurrent variant (RNN) that captures non-Markovian dynamics by allowing updates to depend on trial history. Applied to a large behavioral dataset of mice learning a sensory decision-making task over multiple weeks, our models improved predictions on held-out data. The inferred rules revealed asymmetric updates after correct versus error trials and history dependence, consistent with non-Markovian learning. Overall, these results introduce a flexible framework for inferring biological learning rules from behavioral data in de novo learning tasks—providing insights to inform experimental training protocols and the development of behavioral digital twins.
@@ 1 Introduction
While inference of learning rules is gaining traction, prior studies have key limitations. For example, many leverage reinforcement learning (RL) paradigms with normative approaches that assume reward maximization [15, 16, 17] or weight update under a parametric learning rule (e.g., Q-learning [18], REINFORCE [19]). Although these approaches have provided clear insights into animal learning, they may lack the flexibility to capture the detailed structure of real learning rules, which are often unknown and may not align with parametric forms conceived by human researchers. As a result, existing models may span only a small subset of the space of possible learning strategies animals actually use. Recent work has proposed more flexible quantitative models of learning based on recurrent neural networks (RNNs) [20, 21, 22, 23] or mechanisms proposed by large language models [24]. However, these studies have generally focused on “bandit” style tasks, in which animals adapt their behavior to time-varying rewards in a fixed task structure, as opposed to learning a novel task or behavior from scratch [25, 26]. De novo learning is common in experimental neuroscience, as animals are routinely trained to acquire novel input–output mappings from scratch, yet comparatively understudied. Finally, a variety of descriptive models have been developed to characterize trial-by-trial changes in weights that govern an animal’s policy as it evolves over learning, but do not identify any rules governing these weight updates [27, 28].
@@
    Insights into animal learning strategies. We analyze the inferred learning rules and identify several departures from classic policy gradient learning, including negative “baseline” in updates, decision outcome asymmetry, and non-Markovian reward history dependencies lasting multiple trials (Fig. 5).
@@ 2 Model and methods
Model of decision-making. We begin with a dynamic Bernoulli generalized linear model (GLM) of an animal’s time-varying behavior during learning [27, 29, 19]. This provides an interpretable model of an animal’s decision-making policy in terms of a set of weights that evolve dynamically over the course of learning. This differs from recent work in which the animal’s policy is directly parametrized by an RNN [20, 21], which offers increased flexibility but may be challenging to interpret.
@@
We treat the initial weights $\mathbf{w}_{0}$ as trainable, initialized via psychometric curve estimates (Supp. C). The inferred weights $\mathbf{w}_{t}$ parameterize the GLM response distribution (Eq. 1), and network parameters $\theta$ are optimized to maximize the log-likelihood of observed choices:
@@
2:Trainable parameters: Neural network weights and biases; initial GLM weight estimates $\mathbf{w}_{0}$
3:for trial $t=1$ to $T$ do
4:  Construct input vector: $u_{t}=[s_{t},y_{t},r_{t},\mathbf{w}_{t}]$
5:  Compute weight updates: $\Delta\mathbf{w}_{t}\leftarrow\mathrm{NN}(u_{t})$
6:  Update the inferred GLM weights: $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}+\Delta\mathbf{w}_{t}$
7:  Predict choice probability: $P_{y_{t}}=\sigma(\mathbf{w}_{t}^{\top}x_{t})$
8:  Accumulate binary cross-entropy loss: $\mathcal{L}_{t}=-\left[y_{t}\log P_{y_{t}}+(1-y_{t})\log(1-P_{y_{t}})\right]$
9:end for
10:Objective: Minimize total loss $\mathcal{L}=\sum_{t=1}^{T}\mathcal{L}_{t}$ using gradient descent
@@ 3 Simulation results: recovery of ground-truth learning rules
Figure 3: Simulated example with a non-Markovian learning rule. The RNNGLM recovered the reward history dependence present in the ground truth: larger weight updates occur when the past $3$ trials were rewarded (black) versus unrewarded (red). Such differences were absent in DNNGLM, as it can only account for Markovian learning rules, where weight updates depend on current trial information. The ability of RNNGLM to capture these effects reflects its greater flexibility. Minor deviations from the ground truth are likely due to finite data (Fig. 2C). To keep the figure simple, we only plotted with the learning rule given stimulus weight $w_{stim}=0$ and after correct choices, but we observed similar phenomena for other parameter settings.
@@
We also add more results in the Supp. Material demonstrating that our main findings are robust across a variety of conditions: (i) a learning rule without the decision outcome assymetry (Supp.  Fig. 6); (ii) recovery of an in-between learning rule when mixing update functions across individuals (Supp.  Fig. 11); (iii) different initial stimulus weights (Supp.  Fig. 9); (iv) longer trial lengths (Supp.  Fig. 10). We also show sensitivity to the initial weight estimate (Supp.  Table 3) and degradation of recovery under added noise and incorrect initial weight estimates (Supp.  Table 4 and 5).
@@ 4 Application to mouse learning data
Figure 5: Properties of inferred learning rules. (A) For the dataset in Fig. 4, we observe the following deviations from the REINFORCE learning rule: (i) DNNGLM (left) and RNNGLM (middle) both reveal a suboptimal negative baseline (where weight updates are negative on incorrect trials, even though they should be positive to learn this task), (ii) a side asymmetry (distinct updates for left $s<0$ versus right side $s>0$), and (iii) an asymmetry in updates after correct (solid) versus incorrect (dashed) decisions. Bias weight plot is provided in Supp. Fig. 8; note $\mathbf{w}$ is fixed at positive weights because the training data mainly involved positive $\mathbf{w}$. We follow the same plotting convention in Fig. 2; since RNNGLM depends on additional historical variables, we plot the mean $\Delta w_{stim}$ averaged across these historical dimensions. (B) RNNGLM suggests non-Markovity (history dependence) in learning: larger weight updates are observed when all past trials were rewarded (black) versus unrewarded (red). This gap widens as we condition on history beyond just the previous trial, suggesting that the history dependency extends beyond the most recent past trial. We note that plotting up to four past trials was intended only as an illustrative example, not a fundamental limit of RNNGLM; as shown in Appendix Fig. 13, the model suggests longer history dependencies. To keep the figure simple, we only plotted with stimulus $w$ fixed at 0, but similar trends are observed elsewhere.
@@
Moreover, our methods revealed empirical features of mouse learning that deviate from classical learning rules (Fig. 5A). These include the presence of a “negative baseline,” where weight updates following incorrect trials actually decreased task accuracy — even though a positive update would aid learning in this particular task. We also observed asymmetries in update magnitude based on decision correctness, with distinct updates following correct (solid lines) versus incorrect (dashed lines) choices, as well as asymmetries across stimulus sides (i.e., different update functions for left vs. right choices). Notably, such side asymmetries have also been reported in prior work [33].
@@
Because inferring flexible learning rules from binary choices is an under-constrained inverse problem, multiple parameterizations can fit the data similarly well. We therefore repeated inference for each model (DNNGLM, RNNGLM) across multiple random seeds, holding data splits and hyperparameters fixed. Across seeds, log-likelihoods were similar, while the recovered update functions showed modest variation. To visualize this, we now plot mean and standard deviation across of weight updates across seeds (Appendix Fig. 12). Despite some variability, the key characteristics are consistent across seeds—a negative baseline, side asymmetry, and multi-trial (non-Markovian) dependence on reward/stimulus history. Moreover, "ablations" suggest these features are functionally meaningful: enforcing a non-negativity constraint on the baseline reduces held-out predictive performance (Table 1). Additionally, to assess fairness, we matched parameter counts between DNNGLM and RNNGLM, and observed that RNNGLM still outperformed ( p=5.53e4p=5.53e-4), indicating the gap is not due to just model size.
Finally, Proposition 1 in Supp. A provides theoretical support for our approach in highly simplified settings. A full investigation of identifiability and convergence is left for future work.
@@
Table 1: Animal-level held-out log-likelihoods (higher is better) for Fig. 4B, except that we report total log-likelihood (summed across animals) instead of relative log-likelihood. We show $p$-values from paired $t$-tests across mice, comparing different learning rule models on the IBL dataset. Each entry shows two values (one per model in the comparison), reported as mean ± standard error across cross-validation seeds. For the interpretable RNN baseline — which predicts behavior directly rather than parameterizing weights or weight updates — we used tinyRNN [20] with 8 GRUs, more than the 1–4 GRUs used in their original paper, which performed even worse in our tests. Interestingly, removing the suboptimal negative baseline from REINFORCE leads to significantly worse performance. Results for held-out future data also show consistent trends (Supp.  Table 2).
@@ 5 Related Work
Methods for inferring learning rules during the acquisition of new task structures remain sparse. Those that do exist typically assume a predefined functional form [18, 19, 33], limiting their ability to capture learning strategies that differ from hand-designed models. In addition, many are memoryless or ’Markovian’, assuming that weight updates depend only on the task variables encountered on the current trial; this stands in contrast to biological learning, where history plays a prominent role — such as through eligibility traces [52, 53] or the influence of trial history on decision-making [29, 54, 55, 56]. In this work, we focus on inferring learning rules from behavioral data, motivated by the broad availability of behavioral datasets and the potential for large-scale inference without requiring simultaneous neural recordings. In parallel, several studies have proposed approaches to infer synaptic-level plasticity rules [57, 58, 59, 60, 61, 62, 63, 64, 65], which offer detailed insights into neural mechanisms. While some of these approaches can also be applied to behavioral data (e.g., [61]), their focus has largely been on neural-level inference and often requires neural activity recordings or handcrafted functional bases for interpretation. Our method instead emphasizes flexible inference directly from behavior, improving weight recovery via psychometric-curve initialization and extending flexibility with recurrent networks to capture multi-trial history beyond bandit tasks.
@@ 6 Discussion
Broader outlook for neuroscience and AI. By inferring learning rules directly from behavior without assuming a fixed model structure, our approach offers a data-driven framework for understanding how animals learn de novo. This has several implications. First, interpreting the inferred learning rules could offer insights into learning behavior (e.g., suboptimal or history-dependent patterns), which may point to additional variables—such as trial history, internal states, or reward timing—that are worth tracking in future behavioral learning studies. In turn, more accurate behavioral learning models could inform applications in health and conservation—e.g., anticipating how animals respond to environmental changes or interventions [4, 5, 6, 7, 8, 9, 10, 11, 12]. This also lays a foundation for more challenging settings with richer task structures and scaling to more expressive architectures. Second, de novo learning is ubiquitous in experimental neuroscience, where animals are often trained from scratch on novel tasks, yet remains understudied. By revealing structure in how animals learn during this phase, our method can inform experimental design and training protocols, supporting more efficient task training [19].
@@
Limitations and future directions. Pooling across animals). Neural network–based approaches are often data hungry, and reconstruction accuracy improves with more data (Fig. 2C). To address this, we pool across animals rather than fitting individual subjects. While pooling is common in neuroscience [85, 86, 87], it can obscure individual variability and inherit known limitations [88]. Nonetheless, learning rule inference from behavior alone is underconstrained in low-data regimes, and pooling offers a practical mitigation strategy. This population-level approach aligns with broader trends in neuroscience and AI toward large-scale behavioral datasets.
@@
Static vs. dynamic learning rules). Our current model assumes a static weight update function. While RNNs may implicitly capture temporal changes—perhaps explaining why fitting separate RNNGLMs to the first and second halves of training did not yield improvement—future work could explicitly model dynamic learning rules to better disentangle non-Markovian dependencies from nonstationarity. An important direction is to investigate whether the observed suboptimal negative baseline arises from specific cognitive mechanisms — such as choice perseverance or forgetting [89] — or from model mismatch. More broadly, suboptimal learning is common in the literature [90, 91], especially in pathological populations [92, 93], where applying our approach could be interesting.
@@
Stochasticity). Our current learning model is deterministic. Capturing stochasticity in learning would be a valuable extension.
@@
Generalization across datasets, tasks, and architectures). We evaluated our method using the same preprocessed IBL dataset as [19]. This focus on a relatively simple task was intentional: it mirrors widely used de novo learning paradigms in laboratory neuroscience, where animals are introduced to novel tasks with minimal prior structure. Such simplified tasks capture key hallmarks of biological learning—history dependence, biases, and suboptimal strategies—while providing a tractable platform for reverse-engineering learning dynamics. In this sense, our work should be viewed as a stepping stone toward broader generalization: establishing validity in a controlled setting before extending to richer tasks and datasets. Indeed, to demonstrate potential scalability, we tested higher-dimensional inputs and found that our approach continues to recover core qualitative features of learning rules, with performance improving as data availability increases (Appendix Table 6). These results indicate that our framework can extend beyond the simple tasks considered here.
@@
Beyond extending to diverse tasks and datasets, another future avenue would be to investigate how different architectures—and their associated inductive biases—affect inference accuracy, particularly under data limitations. Our framework currently fixes the decision model (standard models like GLM): this aids interpretability and follows common neuroscience practice, but limits expressivity. We view this modular approach—using a standard decision model while focusing on trial-by-trial weight updates—as a principled first step, consistent with strategies in neuroscience and machine learning where components are studied separately before joint inference. Extending the framework to jointly infer both the decision model and the learning rule is an important direction for future work.

### v2

# Flexible inference for animal learning rules using neural networks

@@ Abstract
Understanding how animals learn is a central challenge in neuroscience, with growing relevance to the development of animal- or human-aligned artificial intelligence. However, existing approaches tend to assume fixed parametric forms for the learning rule (e.g., Q-learning, policy gradient), which may not accurately describe the complex forms of learning employed by animals in realistic settings. Here we address this gap by developing a framework to infer learning rules directly from behavioral data collected during de novo task learning. We assume that animals follow a decision policy parameterized by a generalized linear model (GLM), and we model their learning rule—the mapping from task covariates to per-trial weight updates—using a deep neural network (DNN). This formulation allows flexible, data-driven inference of learning rules while maintaining an interpretable form of the decision policy itself. To capture more complex learning dynamics, we introduce a recurrent neural network (RNN) variant that relaxes the Markovian assumption that learning depends solely on covariates of the current trial, allowing for learning rules that integrate information over multiple trials. Simulations demonstrate that the framework can recover ground-truth learning rules. We applied our DNN and RNN-based methods to a large behavioral dataset from mice learning to perform a sensory decision-making task and found that they outperformed traditional RL learning rules at predicting the learning trajectories of held-out mice. The inferred learning rules exhibited reward-history–dependent learning dynamics, with larger updates following sequences of rewarded trials. Overall, these methods provide a flexible framework for inferring learning rules from behavioral data in de novo learning tasks, setting the stage for improved animal training protocols and the development of behavioral digital twins.
@@ 1 Introduction
While inference of learning rules is gaining traction, prior studies have key limitations. For example, many leverage reinforcement learning (RL) paradigms with normative approaches that assume reward maximization [24, 78, 98] or weight update under a parametric learning rule (e.g., Q-learning [59], REINFORCE [5, 35]). Although these approaches have provided clear insights into animal learning, they may lack the flexibility to capture the detailed structure of real learning rules, which are often unknown and may not align with parametric forms conceived by human researchers. As a result, existing models may span only a small subset of the space of possible learning strategies animals actually use. Recent work has proposed more flexible quantitative models of learning based on recurrent neural networks (RNNs) [52, 72, 29, 25] or mechanisms proposed by large language models [15]. However, these studies have generally focused on “bandit” style tasks, in which animals adapt their behavior to time-varying rewards in a fixed task structure, as opposed to learning a novel task or behavior from scratch [16, 41]. De novo learning is common in experimental neuroscience, as animals are routinely trained to acquire novel input–output mappings from scratch, yet comparatively understudied. Finally, a variety of descriptive models have been developed to characterize trial-by-trial changes in weights that govern an animal’s policy as it evolves over learning, but do not identify any rules governing these weight updates [93, 12].
@@
    Insights into animal learning strategies. We analyze the inferred learning rules and identify several departures from classic policy gradient learning, including non-Markovian reward history dependencies lasting multiple trials (Fig. 5).
@@ 2 Model and methods
Model of decision-making. We begin with a dynamic Bernoulli generalized linear model (GLM) of an animal’s time-varying behavior during learning [93, 92, 5, 35]. This provides an interpretable model of an animal’s decision-making policy in terms of a set of weights that evolve dynamically over the course of learning. This differs from recent work in which the animal’s policy is directly parametrized by an RNN [52, 72], which offers increased flexibility but may be challenging to interpret.
@@
Initial weights $\mathbf{w}_{0}$ can be treated as trainable parameters, initialized via psychometric curve estimates (Supp. C). The inferred weights $\mathbf{w}_{t}$ parameterize the GLM response distribution (Eq. 1), and network parameters $\theta$ are optimized to maximize the log-likelihood of observed choices:
@@
2:for trial $t=1$ to $T$ do
3:  Construct input vector: $u_{t}=[s_{t},y_{t},r_{t},\mathbf{w}_{t}]$
4:  Compute weight updates: $\Delta\mathbf{w}_{t}\leftarrow\mathrm{NN}(u_{t})$
5:  Update the inferred GLM weights: $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}+\Delta\mathbf{w}_{t}$
6:  Predict choice probability: $P_{y_{t}}=\sigma(\mathbf{w}_{t}^{\top}x_{t})$
7:  Accumulate binary cross-entropy loss: $\mathcal{L}_{t}=-\left[y_{t}\log P_{y_{t}}+(1-y_{t})\log(1-P_{y_{t}})\right]$
8:end for
9:Objective: Minimize total loss $\mathcal{L}=\sum_{t=1}^{T}\mathcal{L}_{t}$ using gradient descent
@@ 3 Simulation results: recovery of ground-truth learning rules
Figure 3: Simulated example with a non-Markovian learning rule. The RNNGLM recovered the reward history dependence present in the ground truth: larger weight updates occur when the past $3$ trials were rewarded (black) versus unrewarded (red). Such differences were absent in DNNGLM, as it can only account for Markovian learning rules, where weight updates depend on current trial information. The ability of RNNGLM to capture these effects reflects its greater flexibility. Minor deviations from the ground truth are likely due to finite data effects (Fig. 2C). To keep the figure simple, we only plotted with the learning rule given stimulus weight $w_{stim}=0$ and after correct choices, but we observed similar phenomena for other parameter settings.
@@
We also add more results in the Supp. Material demonstrating that our main findings are robust across a variety of conditions: (i) a learning rule without the decision outcome asymmetry (Supp. Fig. 6); (ii) recovery of an in-between learning rule when mixing update functions across individuals (Supp. Fig. 11); (iii) different initial stimulus weights (Supp. Fig. 9); (iv) longer trial lengths (Supp. Fig. 10). While the initial policy weights are typically unknown, we estimated them from the psychometric curve (Supp. C) and demonstrated recovery in Supplementary Fig. 10. For faster simulations in the main text, and to allow multiple repetitions for robustness testing, we used shorter trial sequences. These shorter sequences make it difficult to estimate the initial weight due to fewer samples at the beginning, so we initialized with a known initial weight in this case. Robustness tests showing recovery with incorrect $W_{0}$ are provided in Supplementary Table 5. We also show sensitivity to the initial weight estimate (Supp. Table 3) and degradation of reconstruction under added noise (Supp. Table 4).
@@ 4 Application to mouse learning data
Figure 5: Properties of inferred learning rules. (A) For the dataset in Fig. 4, DNNGLM reveals negative weight updates following errors (coined as ‘negative baseline’ in [35], which was also observed in our flexible framework). Bias weight plot is provided in Supp. Fig. 8; note $\mathbf{w}$ is fixed at positive weights because the training data mainly involved positive $\mathbf{w}$. We follow the same plotting convention in Fig. 2; since RNNGLM depends on additional historical variables, we plot the mean $\Delta w_{stim}$ averaged across these historical dimensions. (B) RNNGLM suggests non-Markovianity (history dependence) in learning: larger weight updates are observed when all past trials were rewarded (black) versus unrewarded (red). This gap widens as we condition on history beyond just the previous trial, suggesting that the history dependency extends beyond the most recent past trial. We note that plotting up to four past trials was intended only as an illustrative example, not a fundamental limit of RNNGLM; as shown in Supp. Fig. 13, the model suggests longer history dependencies. To keep the figure simple, we only plotted with stimulus $w$ fixed at 0, but similar trends are observed elsewhere.
@@
Moreover, our methods revealed empirical features of mouse learning that deviate from classical learning rules (Fig. 5A). These include the presence of a “negative baseline” (coined in [35] but also observed in our flexible framework), where weight updates following incorrect trials actually decreased task accuracy — even though a positive update would aid learning in this particular task. We also observed asymmetries in update magnitude based on decision correctness, with distinct updates following correct (solid lines) versus incorrect (dashed lines) choices, as well as asymmetries across stimulus sides (i.e., different update functions for left vs. right choices). Notably, such side asymmetries have also been reported in prior work [60].
@@
Because inferring flexible learning rules from binary choices is an under-constrained inverse problem, multiple parameterizations can fit the data similarly well. We therefore repeated inference for each model (DNNGLM, RNNGLM) across multiple random seeds, holding data splits and hyperparameters fixed. Across seeds, log-likelihoods were similar, while the recovered update functions showed modest variation. To visualize this, we now plot mean and standard deviation across weight updates across seeds (Supp. Fig. 12). Despite some variability, the key characteristics are consistent across seeds—a negative baseline, side asymmetry, and multi-trial (non-Markovian) dependence on reward/stimulus history. Moreover, the elimination of these features suggests they are functionally meaningful: enforcing a non-negativity constraint on the baseline reduces held-out predictive performance (Table 1). Additionally, to assess fairness, we matched parameter counts between DNNGLM and RNNGLM, and observed that RNNGLM still outperformed (p=5.53e4p=5.53e-4), indicating the gap is not due to just model size.
@@
Table 1: Animal-level held-out log-likelihoods (higher is better) for Fig. 4B, except that we report total log-likelihood (summed across animals) instead of relative log-likelihood. We show $p$-values from paired $t$-tests across mice, comparing different learning rule models on the IBL dataset. Each entry shows two values (one per model in the comparison), reported as mean ± standard error across cross-validation seeds. For the interpretable RNN baseline — which predicts behavior directly rather than parameterizing weights or weight updates — we used tinyRNN [52] with 8 GRUs, more than the 1–4 GRUs used in their original paper, which performed even worse in our tests. Results for held-out future data also show consistent trends (Supp. Table 2).
@@ 5 Related Work
Methods for inferring learning rules during the acquisition of new task structures remain sparse. Those that do exist typically assume a predefined functional form [59, 5, 60, 35], limiting their ability to capture learning strategies that differ from hand-designed models. In addition, many are memoryless or ’Markovian’, assuming that weight updates depend only on the task variables encountered on the current trial; this stands in contrast to biological learning, where history plays a prominent role — such as through eligibility traces [67, 39] or the influence of trial history on decision-making [92, 100, 75, 1]. In this work, we focus on inferring learning rules from behavioral data, motivated by the broad availability of behavioral datasets and the potential for large-scale inference without requiring simultaneous neural recordings. In parallel, many studies have proposed approaches to study synaptic-level plasticity rules [62, 86, 55, 76, 70, 20, 19, 88, 9, 63, 64, 61, 45]. While some of these approaches can also be applied to behavioral data (e.g., [70]), their focus has largely been on synaptic-level inference and often requires neural activity recordings or handcrafted functional bases for interpretation. Our method instead emphasizes flexible inference directly from behavior, improving weight recovery via psychometric-curve initialization and extending flexibility with recurrent networks to capture multi-trial history beyond bandit tasks.
@@ 6 Discussion
Broader outlook for neuroscience and AI. By inferring learning rules directly from behavior without assuming a fixed model structure, our approach offers a data-driven framework for understanding how animals learn de novo. This has several implications. First, interpreting the inferred learning rules could offer insights into learning behavior (e.g., suboptimal or history-dependent patterns), which may point to additional variables—such as trial history, internal states, or reward timing—that are worth tracking in future behavioral learning studies. In turn, more accurate behavioral learning models could inform applications in health and conservation—e.g., anticipating how animals respond to environmental changes or interventions [40, 94, 68, 42, 105, 47, 27, 84, 77]. This also lays a foundation for more challenging settings with richer task structures and scaling to more expressive architectures. Second, de novo learning is ubiquitous in experimental neuroscience, where animals are often trained from scratch on novel tasks, yet remains understudied. By revealing structure in how animals learn during this phase, our method can inform experimental design and training protocols, supporting more efficient task training [5, 35].
@@
Limitations and future directions. Pooling across animals. Neural network–based approaches are often data hungry, and reconstruction accuracy improves with more data (Fig. 2C). To address this, we pool across animals rather than fitting individual subjects. While pooling is common in neuroscience [4, 3, 10], it can obscure individual variability and inherit known limitations [74]. Nonetheless, learning rule inference from behavior alone is underconstrained in low-data regimes, and pooling offers a practical mitigation strategy. This population-level approach aligns with broader trends in neuroscience and AI toward large-scale behavioral datasets.
@@
Static vs. dynamic learning rules. Our current model assumes a static weight update function. While RNNs may implicitly capture temporal changes—perhaps explaining why fitting separate RNNGLMs to the first and second halves of training did not yield improvement—future work could explicitly model dynamic learning rules to better disentangle non-Markovian dependencies from nonstationarity. An important direction is to investigate whether the observed suboptimal negative baseline arises from specific cognitive mechanisms — such as choice perseverance or forgetting [23] — or from model mismatch. More broadly, suboptimal learning is common in the literature [73, 66], especially in pathological populations [26, 56], where applying our approach could be interesting.
@@
Stochasticity. Our current learning model is deterministic. Capturing stochasticity in learning would be a valuable extension.
@@
Generalization across datasets, tasks, and architectures. We evaluated our method using the same preprocessed IBL dataset as [5]. This focus on a relatively simple task was intentional: it mirrors widely used de novo learning paradigms in laboratory neuroscience, where animals are introduced to novel tasks with minimal prior structure. Such simplified tasks capture key hallmarks of biological learning—history dependence, biases, and suboptimal strategies—while providing a tractable platform for reverse-engineering learning dynamics. In this sense, our work should be viewed as a stepping stone toward broader generalization: establishing validity in a controlled setting before extending to richer tasks and datasets. Indeed, to demonstrate potential scalability, we tested higher-dimensional inputs and found that our approach continues to recover core qualitative features of learning rules, with performance improving as data availability increases (Supp. Table 6). These results indicate that our framework can extend beyond the simple tasks considered here.
@@
Beyond extending to diverse tasks and datasets, another future avenue would be to investigate how different architectures—and their associated inductive biases—affect inference accuracy, particularly under data limitations. Our framework currently fixes the decision model (e.g., standard models such as GLMs), which aids interpretability and follows common neuroscience practice but limits expressivity. In fact, for multiplicative decision processes, this assumption leads to a significant performance degradation (Supp. Table 7). We view this modular approach—using a standard decision model while focusing on trial-by-trial weight updates—as a principled first step, consistent with strategies in neuroscience and machine learning where components are studied separately before joint inference. Extending the framework to jointly infer both the decision model and the learning rule is an important direction for future work.